40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 27 20:47:00 2007, tobin, Update, PSL, FSS fss-tf0001.pdf
    Reply  Wed Nov 28 04:21:46 2007, rana, Update, PSL, FSS 
Message ID: 128     Entry time: Wed Nov 28 04:21:46 2007     In reply to: 127
Author: rana 
Type: Update 
Category: PSL 
Subject: FSS 

Quote:
Rana, Tobin

We looked at the RF PD signal to the FSS (siphoning off a signal via a minicircuits directional coupler) and also took an open loop transfer function of the FSS. In the transfer function we saw the step at 100 kHz (mentioned by Rob) as well as some peculiar behavior at high frequency. The high frequency behavior (with a coupling of ~ -20 dB) turns out to be bogus, as it is still present even with the beam blocked. Rearranging the cabling had no effect; the cause is apparently inside the FSS. The step at 100 kHz turns out to be a saturation effect, as it moved as we lowered the signal amplitude, disappearing as we approached -60 dBm. (Above the step, the measurement data is valid; below, bogus.)

Transfer functions will be attached to this entry.

Some things to check tomorrow: the RF signal to the PC, RF AM generation by the PC, LO drive level into the FSS, RF reflection from the PC, efficiency of FSS optical path, quality of RF cabling.


I would also add to Tobin's entry that we believe what Rob was seeing was saturation.

With the bi-directional coupler in there, the RF signal into the FSS board clearly went UP if moved the offset slider away from zero.
With a scope looking at the IN2 testpoint, we can see that there's less than 2 mV offset at zero slider offset.

One tangential thing we noticed with the coupler is that, in lock, the amount of reflected RF is around the same as that going in to the mixer.
I have always wanted to look at this but have only had uni-directional couplers in the past. I think that the double balanced mixer is inherently
not a 50 Ohm device during the times where the diodes are being switched. IF that's the case we might do better in the future by having an RF
buffer on board just before the mixer to isolate the PD head from these reflections.
ELOG V3.1.3-